OpenStudy (anonymous):

Diff >> tan^-1(8/w) =??

hartnn (hartnn):

u know chain rule ?

OpenStudy (zepp):

U u substitution to solve!

OpenStudy (zepp):

Loi, use chain rule, not sawaken yet ;(

hartnn (hartnn):

what is d/dx (tan^-1 x) =?

OpenStudy (anonymous):

answer is 1/x^2 +1

hartnn (hartnn):

yes, so what will be d/dw (tan^-1 8/w) ?

hartnn (hartnn):

just replace x by 8/w

OpenStudy (anonymous):

8/w^2 + 64

hartnn (hartnn):

nopes. it will be \(\huge \frac{1}{(8/w)^2+1}.\frac{d}{dw}{(8/w)}\) this was using chain rule, did u understand this ?

hartnn (hartnn):

now can u differentiate 8/w ??

OpenStudy (anonymous):

d/dw (8/w) = -8/w^2

hartnn (hartnn):

good :) so u have \(\huge \frac{1}{(8/w)^2+1}.\frac{-8}{w^2}\) can u simplify this?

OpenStudy (anonymous):

\[\frac{ d }{ aw } \tan ^{-1}(\frac{ 8 }{ w }) = \frac{ \frac{ -8 }{ e } }{ 1+\frac{ 64 }{ w ^{2} } }\]

OpenStudy (anonymous):

yes or no ??

hartnn (hartnn):

i would rather simplify it like this : \(\huge \frac{1}{64/w^2+1}.\frac{-8}{w^2}=\frac{-8}{w^2+64}\) got this ?

OpenStudy (anonymous):

Thank you very much.

hartnn (hartnn):

Welcome very much ^_^

OpenStudy (anonymous):

:)

hartnn (hartnn):

any more questions/doubts?

OpenStudy (anonymous):

No questions thank you. :)